Sinclair Manufacturing and Boswell Brothers Inc. are both involved in the production of tile for the home-building industry. Their financial information is as follows:

Capital Structure

Sinclair Boswell
Debt @12% 600,000 0
common stock, $10 per share 400,000 1,000,000
$1,000,000 $1,000,000
common shares 40,000 100,000
Operating Plan
Sales (50,000 units at $20 each) $1,000,000 $1,000,000
Less: Variable costs 800,000 500,000
($ 16 per unit) ($10 per unit)
Fixed costs 0 300,000
Earnings before interest and taxes (EBIT). $ 200,000 $200,000

a. If you combine Sinclair's capital structure with Boswell's operating plan, what is the DCL?

b. If you combine Boswell's capital structure with Sinclair's operating plan, what is the DCL?

c. Explain why you got the results you did in parts a and b.

d. In part b, if sales double, by what percentage will ESP increase?
Answer #1

DCL=DEGREE OF COMBINED LEVERAGE

VC= VARIABLE COST

FC= FIXED COST

I=INTEREST AMOUNT

Q= QUANTITY

P= PRICE= SELLING PRICE

a) Interest Expense (debt*rate)= $600,000 x 12% = 72,000

DCL = Q(P-VC)/[Q(P-VC)-FC-I]

= 50,000(20-10)/[50,000(20-10) - 300,000 - 72,000]

= 500,000/128000

= 3.91

b)

DCL = Q(P-VC)/[Q(P-VC)-FC-I]

= 50,000(20-10)/[50,000(20-10) - 0 - 0]

= 500,000/500,000 = 1

=

c. in part a interest and fixed cost is added but in part b, interest and fixed cost is zero so result is different

EPS will increase by 100 percent.There is no leverage involved, EPS merely grows at the same rate as sales.